# NeuroBOLT: Resting-state EEG-to-fMRI Synthesis with Multi-dimensional Feature Mapping

### v2

@@ 3.3 Intra-subject Prediction
Intra-scan BIOT[59] 0.531$\pm$0.223 0.518$\pm$0.207 0.490$\pm$0.162 0.459$\pm$0.110 0.410$\pm$0.205 0.411$\pm$0.231 0.493$\pm$0.133 0.473
LaBraM[22] 0.540$\pm$0.176 0.519$\pm$0.197 0.493$\pm$0.153 0.490$\pm$0.176 0.411$\pm$0.179 0.449$\pm$0.177 0.487$\pm$0.167 0.484
BEIRA [25] 0.357$\pm$0.241 0.396$\pm$0.240 0.294$\pm$0.228 0.320$\pm$0.220 0.234$\pm$0.194 0.328$\pm$0.197 0.456$\pm$0.240 0.341
Li, et al. [31] 0.460$\pm$0.228 0.515$\pm$0.207 0.376$\pm$0.169 0.457$\pm$0.204 0.324$\pm$0.183 0.398$\pm$0.194 0.583$\pm$0.170 0.445
NeuroBOLT (ours) 0.588$\pm$0.166 0.566$\pm$0.183 0.502$\pm$0.168 0.559$\pm$0.141 0.437$\pm$0.184 0.480$\pm$0.213 0.587$\pm$0.162 0.531
Inter-subject FFCL [29] 0.326$\pm$0.094 0.412$\pm$0.039 0.327$\pm$0.078 0.437$\pm$0.091 0.243$\pm$0.125 0.373$\pm$0.082 0.512$\pm$0.048 0.376
CNN Transformer [44] 0.218$\pm$0.204 0.412$\pm$0.114 0.298$\pm$0.097 0.316$\pm$0.153 0.232$\pm$0.086 0.180$\pm$0.106 0.282$\pm$0.185 0.273
STT Transformer [48] 0.269$\pm$0.197 0.188$\pm$0.056 0.226$\pm$0.130 0.280$\pm$0.143 0.074$\pm$0.126 0.142$\pm$0.101 0.347$\pm$0.124 0.218
BIOT [59] 0.457$\pm$0.123 0.512$\pm$0.039 0.393$\pm$0.128 0.445$\pm$0.084 0.299$\pm$0.063 0.413$\pm$0.073 0.529$\pm$0.110 0.435
LaBraM [22] 0.177$\pm$0.116 0.211$\pm$0.105 0.153$\pm$0.132 0.170$\pm$0.152 0.047$\pm$0.111 0.147$\pm$0.122 0.150$\pm$0.152 0.151
BEIRA [25] 0.421$\pm$0.112 0.482$\pm$0.063 0.384$\pm$0.147 0.452$\pm$0.149 0.241$\pm$0.135 0.410$\pm$0.097 0.492$\pm$0.106 0.412
Li, et al. [31] 0.505$\pm$0.063 0.430$\pm$0.048 0.415$\pm$0.114 0.416$\pm$0.076 0.217$\pm$0.139 0.424$\pm$0.072 0.529$\pm$0.092 0.419
NeuroBOLT (ours) 0.482$\pm$0.100 0.561$\pm$0.046 0.423$\pm$0.115 0.496$\pm$0.136 0.335$\pm$0.144 0.453$\pm$0.106 0.564$\pm$0.115 0.473
@@ 3.4 Inter-subject Prediction
To assess the generalizability of NeuroBOLT across subjects, we first evaluated its performance on held-out recordings in the resting-state dataset, this time using EEG data to predict the fMRI ROI time series across entire scans (with length of about 20 minutes). As shown in the lower part of Table 1, our model achieves the highest average correlation across ROIs, with a score of 0.473. Figure 5 shows the correspondence between the correlation values and the appearance of the reconstructions. This visualization reveals that NeuroBOLT successfully captures major features of fMRI dynamics in reconstructing the whole unseen fMRI scan, especially for primary sensory regions and global signals. Thus, even though both EEG and resting-state fMRI have relatively low signal-to-noise ratio, NeuroBOLT still learns meaningful EEG representations and EEG-fMRI correlations that are crucial for EEG-to-fMRI translation. Additional results including performance distributions with statistical significance, MSE table and examples visualizing the best and the worst fMRI scan reconstructions, are provided in Appendix F for comprehensively displaying the performance. Furthermore, in comparison to other state-of-the-art EEG-fMRI translation approaches [31, 25] and EEG encoding models [59, 22, 29, 44, 48], NeuroBOLT achieves lower mean squared error (MSE) and higher correlation, with at least an 8.74% improvement in average correlations. Similar to the scenario of within-subject prediction, NeuroBOLT shows the highest performance in predicting signals from global and primary sensory regions, achieving average correlations of 0.564 and 0.522, respectively, followed by its performance in high-level cognitive and subcortical regions.
@@ 4 Discussion and Conclusion
In this study, we propose NeuroBOLT, a versatile deep-learning solution for projecting scalp EEG to BOLD fMRI signals. By learning a multi-dimensional EEG representation from spatial, temporal, and spectral domains, NeuroBOLT shows a strong capability to reconstruct resting-state fMRI signals from EEG alone. In addition to subject-specific models, it can also predict fMRI time series from held-out subjects across entire 20-minute scans, which is a first in this field. It is important to note that correlations of 0.4-0.5 are considered strong for predicting moment-by-moment signal fluctuations in functional neuroimaging data, given the low signal-to-noise ratio of EEG/fMRI. It is also worth mentioning that our model reconstructs resting-state fMRI signals from deep subcortical regions like the thalamus (with accuracy comparable to, or exceeding, certain cortical ROIs), using raw EEG data with a relatively small set of electrodes (26), enhancing the capabilities of EEG to map subcortical dynamics. This number of channels is typically insufficient to capture detailed neural activity from such regions using methods like EEG source localization [47]. Moreover, NeuroBOLT supports an arbitrary number of EEG input channels, which allows for flexible application across different EEG setups and across different experimental and clinical settings. Overall, by advancing the field of EEG-fMRI translation, this work may ultimately open new possibilities for non-invasive brain research and cost-effective clinical diagnostics.
@@ Appendix A Notation Table
$TE$ the temporal embedding set in spatiotemporal module
$SE$ the spatial embedding set in spatiotemporal module
@@
$FE$ the frequency embedding set in spectral module
$WE$ the window embedding set in spectral module
@@ Appendix B Related Work
Before the adoption of deep learning methods, research into the correlations between EEG and fMRI suggested the potential for inferring high-dimensional fMRI features based on only partial information from relatively low-dimensional EEG measures. Most of these studies investigate EEG-fMRI relationships by regressing the power in one or more EEG frequency bands against the fMRI time series, after convolving the former with a BOLD hemodynamic response model [38, 26]. However, there is currently no comprehensive, definitive model that summarizes the complex translation from EEG to fMRI, possibly owing to complexities such as a potential state-dependence of neurovascular coupling [51], and research into neural-BOLD relationships is still ongoing [21, 19].
Early pioneering work by [39] applied ridge regression on EEG temporal-spectral features to reconstruct fMRI signals from the visual cortex during an eyes-open-eyes-closed task, and to reconstruct fMRI signals from the amygdala during auditory neurofeedback sessions. Since this study attempted to reconstruct fMRI signals only from task-specific brain regions, the ability of EEG to infer fMRI signals from other regions needs further investigation. Using deep neural networks, [33, 34, 35, 5, 3, 25, 31] explored the capability of encoder-decoder based frameworks for EEG-fMRI translation. Specifically, models developed by [33, 34, 35, 5] mainly focus on reconstructing fMRI volumes, but their effectiveness in recovering fMRI temporal characteristics has yet to be quantitatively assessed. The work of [25] and [31] addressed this gap in evaluating reconstruction performance in the temporal domain by employing sequence-to-sequence models to directly reconstruct fMRI time series in deep brain regions from raw EEG, analyzing the temporal correlation between predicted and actual fMRI signals. However, these studies are constrained by small sample sizes and rely on a subject-specific approach, wherein models are solely trained and tested on different sections of the same individual’s scans. Additionally, most of the existing research has focused on EEG-to-fMRI synthesis during task conditions, including cued eye opening/closing [25, 31, 34, 35, 54, 3], leaving the (fully eyes-closed) resting state condition largely unexplored. This gap in the field may stem from the inherently noisier and more random nature of resting-state data [32], which poses challenges for stable and accurate translation between EEG and fMRI. Yet, resting state is a widely used condition due to its simplicity and ease of applicability in patient populations [37]; therefore, there is a crucial need for models that can successfully operate on resting-state data.
@@ Appendix D Dataset and Preprocessing Details
For the EEG data, gradient and ballistocardiogram (BCG) artifacts were reduced using BrainVision Analyzer 2 (Brain Products, Munich, Germany), following the parameters and procedures described in more detail in [40, 17]. Gradient artifact reduction was performed using the average artifact subtraction technique[1], utilizing volume triggers. After correcting for gradient artifacts, the EEG data were downsampled to 250 Hz.
@@
[…]
@@ E.1 Hyperparameter Settings
[…]
@@ E.2 Baselines
  • Li et al. [31]: A Encoder-Decoder CNN that leverages effective spectral representation learning with sinusoidal activation function for Sequence-to-Sequence EEG-fMRI prediction in deep brain regions during eyes-open-eyes-closed (EOEC) task. We adapt this model by only using the encoder part and then using linear projection to make predictions similar to our approach.
  • BEIRA [25]: An interpretable CNN Encoder-Decoder for deep brain region fMRI reconstruction during EOEC task. We also adapt this model by only using the encoder part and then using linear projection to make predictions similar to our approach.
@@
  • LaBraM [22]: A recent EEG foundation model trained on over 2,500 hours of EEG data to learn the spatiotemporal information of EEG signals, which achieved the state-of-the-art downstream classification task performance. We maintain the original model architecture with a 200-dimensional latent embedding and utilize the pretrained weights from this model (LaBraM-base, the only publicly available checkpoint for LaBraM) for fine-tuning on our datasets.
  • BIOT [59]: A transformer-based flexible bio-signal encoder architecture that learns the temporal-spectral information from EEG signals. Since BIOT and LaBraM share similar input processing, and both produce a one-dimensional latent representation for each sample, we set BIOT’s embedding size to match LaBraM and our model ($d_{hiddenembed}=200$) to ensure a fair comparison, keeping other parameters consistent with the original paper.
  • FFCL [29]: A model that combines a CNN model and LSTM model for learning separate representations, which are merged to make the final prediction.
  • ST-Transformer [48]: a transformer-based spatial-temporal feature learning neural network for EEG classification task.
  • CNN Transformer [44]: A transformer convolutional neural network for automated artifact detection in EEG.
For FFCL, ST-Transformer, and CNN Transformer, we retain the original model architectures as used in the baseline experiments in [59]. For detailed implementation, please refer to their GitHub repository [^4] https://github.com/ycq091044/BIOT/tree/main/model.
@@ F.2 MSE Results in Resting-state Data
Intra-scan BIOT[59] 0.271 0.249 0.237 0.287 0.238 0.274 0.249 0.258
LaBraM[22] 0.276 0.249 0.255 0.273 0.236 0.247 0.277 0.259
BEIRA [25] 0.249 0.244 0.260 0.288 0.239 0.241 0.222 0.249
Li, et al. [31] 0.261 0.226 0.275 0.275 0.249 0.229 0.226 0.249
@@
Inter-subject FFCL [29] 0.225 0.212 0.230 0.205 0.250 0.230 0.189 0.220
CNN Transformer [44] 0.261 0.226 0.236 0.223 0.246 0.248 0.282 0.246
STT Transformer [48] 0.240 0.295 0.299 0.244 0.258 0.255 0.245 0.262
BIOT [59] 0.217 0.186 0.220 0.201 0.239 0.220 0.196 0.211
LaBraM [22] 0.247 0.239 0.256 0.246 0.259 0.255 0.246 0.250
BEIRA [25] 0.231 0.194 0.242 0.197 0.250 0.230 0.201 0.221
Li, et al. [31] 0.193 0.218 0.214 0.219 0.263 0.212 0.196 0.216

### v3

@@ 3.3 Intra-subject Prediction
[…]
@@ 3.4 Inter-subject Prediction
To assess the generalizability of NeuroBOLT across subjects, we first evaluated its performance on held-out recordings in the resting-state dataset, this time using EEG data to predict the fMRI ROI time series across entire scans (with length of about 20 minutes). As shown in the lower part of Table 1, our model achieves the highest average correlation across ROIs, with a score of 0.473. Figure 5 shows the correspondence between the correlation values and the appearance of the reconstructions. This visualization reveals that NeuroBOLT successfully captures major features of fMRI dynamics in reconstructing the whole unseen fMRI scan, especially for primary sensory regions and global signals. Thus, even though both EEG and resting-state fMRI have relatively low signal-to-noise ratio, NeuroBOLT still learns meaningful EEG representations and EEG-fMRI correlations that are crucial for EEG-to-fMRI translation. Additional results including performance distributions with statistical significance, MSE table and examples visualizing the best and the worst fMRI scan reconstructions, are provided in Appendix F for comprehensively displaying the performance. Furthermore, in comparison to other state-of-the-art EEG-fMRI translation approaches Li et al. (2024); Kovalev et al. (2022) and EEG encoding models Yang et al. (2023); Jiang et al. (2024); Li et al. (2022); Peh et al. (2022); Song et al. (2021), NeuroBOLT achieves lower mean squared error (MSE) and higher correlation, with at least an 8.74% improvement in average correlations. Similar to the scenario of within-subject prediction, NeuroBOLT shows the highest performance in predicting signals from global and primary sensory regions, achieving average correlations of 0.564 and 0.522, respectively, followed by its performance in high-level cognitive and subcortical regions.
@@ 4 Discussion and Conclusion
In this study, we propose NeuroBOLT, a versatile deep-learning solution for projecting scalp EEG to BOLD fMRI signals. By learning a multi-dimensional EEG representation from spatial, temporal, and spectral domains, NeuroBOLT shows a strong capability to reconstruct resting-state fMRI signals from EEG alone. In addition to subject-specific models, it can also predict fMRI time series from held-out subjects across entire 20-minute scans, which is a first in this field. It is important to note that correlations of 0.4-0.5 are considered strong for predicting moment-by-moment signal fluctuations in functional neuroimaging data, given the low signal-to-noise ratio of EEG/fMRI. It is also worth mentioning that our model reconstructs resting-state fMRI signals from deep subcortical regions like the thalamus (with accuracy comparable to, or exceeding, certain cortical ROIs), using raw EEG data with a relatively small set of electrodes (26), enhancing the capabilities of EEG to map subcortical dynamics. This number of channels is typically insufficient to capture detailed neural activity from such regions using methods like EEG source localization Song et al. (2015). Moreover, NeuroBOLT supports an arbitrary number of EEG input channels, which allows for flexible application across different EEG setups and across different experimental and clinical settings. Overall, by advancing the field of EEG-fMRI translation, this work may ultimately open new possibilities for non-invasive brain research and cost-effective clinical diagnostics.
@@ Appendix A Notation Table
TETE the temporal embedding set in spatiotemporal module
SESE the spatial embedding set in spatiotemporal module
@@
FEFE the frequency embedding set in spectral module
WEWE the window embedding set in spectral module
@@ Appendix B Related Work
Before the adoption of deep learning methods, research into the correlations between EEG and fMRI suggested the potential for inferring high-dimensional fMRI features based on only partial information from relatively low-dimensional EEG measures. Most of these studies investigate EEG-fMRI relationships by regressing the power in one or more EEG frequency bands against the fMRI time series, after convolving the former with a BOLD hemodynamic response model Mantini et al. (2007); Laufs (2008). However, there is currently no comprehensive, definitive model that summarizes the complex translation from EEG to fMRI, possibly owing to complexities such as a potential state-dependence of neurovascular coupling Turner et al. (2020), and research into neural-BOLD relationships is still ongoing Hillman (2014); Heeger and Ress (2002).
Early pioneering work by Meir-Hasson et al. (2014) applied ridge regression on EEG temporal-spectral features to reconstruct fMRI signals from the visual cortex during an eyes-open-eyes-closed task, and to reconstruct fMRI signals from the amygdala during auditory neurofeedback sessions. Since this study attempted to reconstruct fMRI signals only from task-specific brain regions, the ability of EEG to infer fMRI signals from other regions needs further investigation. Using deep neural networks, Liu and Sajda (2019, 2023a, 2023b); Calhas and Henriques (2022); Bricman and Borst (2021); Kovalev et al. (2022); Li et al. (2024) explored the capability of encoder-decoder based frameworks for EEG-fMRI translation. Specifically, models developed by Liu and Sajda (2019, 2023a, 2023b); Calhas and Henriques (2022) mainly focus on reconstructing fMRI volumes, but their effectiveness in recovering fMRI temporal characteristics has yet to be quantitatively assessed. The work of Kovalev et al. (2022) and Li et al. (2024) addressed this gap in evaluating reconstruction performance in the temporal domain by employing sequence-to-sequence models to directly reconstruct fMRI time series in deep brain regions from raw EEG, analyzing the temporal correlation between predicted and actual fMRI signals. However, these studies are constrained by small sample sizes and rely on a subject-specific approach, wherein models are solely trained and tested on different sections of the same individual’s scans. Additionally, most of the existing research has focused on EEG-to-fMRI synthesis during task conditions, including cued eye opening/closing Kovalev et al. (2022); Li et al. (2024); Liu and Sajda (2023a, b); Wei et al. (2020); Bricman and Borst (2021), leaving the (fully eyes-closed) resting state condition largely unexplored. This gap in the field may stem from the inherently noisier and more random nature of resting-state data Liu (2016), which poses challenges for stable and accurate translation between EEG and fMRI. Yet, resting state is a widely used condition due to its simplicity and ease of applicability in patient populations Lv et al. (2018); therefore, there is a crucial need for models that can successfully operate on resting-state data.
@@ Appendix D Dataset and Preprocessing Details
For the EEG data, gradient and ballistocardiogram (BCG) artifacts were reduced using BrainVision Analyzer 2 (Brain Products, Munich, Germany), following the parameters and procedures described in more detail in Moehlman et al. (2019); Goodale et al. (2021). Gradient artifact reduction was performed using the average artifact subtraction techniqueAllen et al. (2000), utilizing volume triggers. After correcting for gradient artifacts, the EEG data were downsampled to 250 Hz.
@@
The EOG, EMG, and ECG channels are excluded from our analysis, resulting in the final channel number of 26 in resting-state and 31 in task-condition. Before sending into the model, the raw EEG data were first resampled to 200Hz for computational efficiency while maintaining the meaningful frequency components in EEG (typically below 100Hz) according to the Nyquist-Shannon sampling theoremNyquist (1928); Shannon (1949), and then normalized by dividing the EEG data values by 100, since the clean EEG data typically fall into the range between -100-100 $\mu V$. In our study, no additional low-pass or high-pass filter is applied to the EEG signal. We also test the model performance in resting-state inter-subject prediction scenario with a high pass filter (0.5Hz) applied to exclude the influence of low-frequency drift. The results with or without filtering are very close, with only a 0.006 absolute difference in average across all regions, which is also non-significant by t-test ($T=0.1525,p=0.8814,\text{two-sided}$). Therefore, we keep the original settings.
@@ E.1 Hyperparameter Settings
For the spatiotemporal module in NeuroBOLT, we leverage the architecture and pre-trained weights from the state-of-the-art EEG foundation model LaBraM Jiang et al. (2024) (pre-trained weights version: LaBraM-base), and then fine-tune it on our dataset. To keep consistent with the setting of the pre-trained model, we also use a token size of 1 sec (200 timestamps) without overlap for this module, resulting in 16 tokens per channel. Since we have EEG data with $C=26$ channels, the final token number would be $16\times 26+1\text{cls token}=417$ before being sent into the transformer, please see other detailed settings of the model architecture in the "fine-tuning session" of Jiang et al. (2024).
For the multi-scale frequency spectral representation learning module, we set the smallest scale to 0.5 seconds, corresponding to 100 points at our sampling rate of 200 Hz. For the multi-scale learning, we use $l_{3}$ throughout our experiments as it achieves the best average performance across other settings, and it includes four window scales: 0.5 seconds, 1 second, 2 seconds, and 4 seconds. The hidden features derived from these four scales are first projected to the same frequency embedding size (equal to the final hidden embedding size: $d_{hiddenembed}=200$) and then projected to the same temporal embedding shape ($d_{time}=32$, which is derived by $\frac{T}{scale0}$, where $T=3200$ is the total input EEG length 3200 timestamps, and as mentioned $scale0=100$). These projections are both achieved by using a single layer of Linear projection. Therefore the shape of the embeddings before the transformer module is $(32\times 26+1\text{cls token},200)$. In this study, we use a linear transformerWang et al. (2020) instead as our encoder to reduce the complexity and improve the speed. A previous study by Yang et al. (2023) proves that the linear and the vanilla transformer yield very similar performance in biosignal representation learning, while the former one takes less time, so we also adopt the transformer with linear complexity in this study, with 8 heads and 4 transformer layers. We obtain the final embedding from each module by an average pooling step. Then the shape of the hidden embeddings from the spatiotemporal and multi-scale spectral modules are summed up, with a final embedding shape $(200,1)$. Other training details are shown in Table 5.
@@ E.2 Baselines
  • Li et al. Li et al. (2024): A Encoder-Decoder CNN that leverages effective spectral representation learning with sinusoidal activation function for Sequence-to-Sequence EEG-fMRI prediction in deep brain regions during eyes-open-eyes-closed (EOEC) task. We adapt this model by only using the encoder part and then using linear projection to make predictions similar to our approach.
  • BEIRA Kovalev et al. (2022): An interpretable CNN Encoder-Decoder for deep brain region fMRI reconstruction during EOEC task. We also adapt this model by only using the encoder part and then using linear projection to make predictions similar to our approach.
@@
  • LaBraM Jiang et al. (2024): A recent EEG foundation model trained on over 2,500 hours of EEG data to learn the spatiotemporal information of EEG signals, which achieved the state-of-the-art downstream classification task performance. We maintain the original model architecture with a 200-dimensional latent embedding and utilize the pretrained weights from this model (LaBraM-base, the only publicly available checkpoint for LaBraM) for fine-tuning on our datasets.
  • BIOT Yang et al. (2023): A transformer-based flexible bio-signal encoder architecture that learns the temporal-spectral information from EEG signals. Since BIOT and LaBraM share similar input processing, and both produce a one-dimensional latent representation for each sample, we set BIOT’s embedding size to match LaBraM and our model ($d_{hiddenembed}=200$) to ensure a fair comparison, keeping other parameters consistent with the original paper.
  • FFCL Li et al. (2022): A model that combines a CNN model and LSTM model for learning separate representations, which are merged to make the final prediction.
  • ST-Transformer Song et al. (2021): a transformer-based spatial-temporal feature learning neural network for EEG classification task.
  • CNN Transformer Peh et al. (2022): A transformer convolutional neural network for automated artifact detection in EEG.
For FFCL, ST-Transformer, and CNN Transformer, we retain the original model architectures as used in the baseline experiments in Yang et al. (2023). For detailed implementation, please refer to their GitHub repository [^4] https://github.com/ycq091044/BIOT/tree/main/model.
@@ F.2 MSE Results in Resting-state Data
Intra-scan BIOTYang et al. (2023) 0.271 0.249 0.237 0.287 0.238 0.274 0.249 0.258
LaBraMJiang et al. (2024) 0.276 0.249 0.255 0.273 0.236 0.247 0.277 0.259
BEIRA Kovalev et al. (2022) 0.249 0.244 0.260 0.288 0.239 0.241 0.222 0.249
Li, et al. Li et al. (2024) 0.261 0.226 0.275 0.275 0.249 0.229 0.226 0.249
@@
Inter-subject FFCL Li et al. (2022) 0.225 0.212 0.230 0.205 0.250 0.230 0.189 0.220
CNN Transformer Peh et al. (2022) 0.261 0.226 0.236 0.223 0.246 0.248 0.282 0.246
STT Transformer Song et al. (2021) 0.240 0.295 0.299 0.244 0.258 0.255 0.245 0.262
BIOT Yang et al. (2023) 0.217 0.186 0.220 0.201 0.239 0.220 0.196 0.211
LaBraM Jiang et al. (2024) 0.247 0.239 0.256 0.246 0.259 0.255 0.246 0.250
BEIRA Kovalev et al. (2022) 0.231 0.194 0.242 0.197 0.250 0.230 0.201 0.221
Li, et al. Li et al. (2024) 0.193 0.218 0.214 0.219 0.263 0.212 0.196 0.216